ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 133, 2003

Ballico, E. Holomorphic vector bundles on open subsets of Stein manifolds

89

Beltrán, A. On a graph associated to invariant conjugacy classes of finite groups

147

Betsakos, D. On separating conformal annuli and Mori’s ring domain

in R^{n} 1

Bollobás, B., G. R. Brightwell and I. Leader The number of 2-SAT functions

45

Brightwell, G. R. See B. Bollobás

Caranti, A. and M. R. Vaughan-Lee Graded Lie algebras of maximal class. V

157

Henke, A. and A. Regev A-codimensions and A-cocharacters  339

Hernández, F. L. and B. Rodríguez-Salinas Lattice-universal Orlicz spaces on probability spaces

9

von Heydebreck, A. Homotopy properties of certain complexes associated to spherical buildings

369

Iglesias, D. and J. C. Marrero Generalized Lie bialgebras and Jacobi structures on Lie groups

285

Itenberg, I. and E. Shustin Viri theorem and topology of real and complex combinatorial hypersurfaces

189

Konovalov, V. N. and D. Leviatan Shape preserving widths of Sobolev-type classes of s-monotone functions on a finite interval

239

Korry, S. Extensions of Meyers–Ziemer results  357

Leader, I. See B. Bollobás

Leviatan, D. See V. N. Konovalov

Levin, M. and W. Lewis Some mapping theorems for extensional dimension

61

Lewis, W. See M. Levin

Lindenstrauss, A. The Hochschild homology of fatpoints  177

Marrero, J. C. See D. Iglesias

Mason, A. W. and A. Schweizer The minimum index of a nonconvergence subgroup of SL_{2} over an arithmetic domain 29 van Mill, J. On the character and π-weight of homogeneous

compacta

321

Nazarov, F., M. Sodin and A. Volberg Local dimension-free estimates for volumes of sublevel sets of analytic functions

269

Peigné, M. On the Patterson–Sullivan measure of some discrete group of isometries

77

Regev, A. See A. Henke Rodríguez-Salinas, B. See F. L. Hernández Schweizer, A. See A. W. Mason Shustin, E. See I. Itenberg Sodin, M. See F. Nazarov Vaughan-Lee, M. R. See A. Caranti

Villani, M. L. A new characterisation of hereditary PI rings  93

Volberg, A. See F. Nazarov